# Connecting Gravity and Quantum Physics: Primordial Black Holes and the Evolution of the Universe

### v2

@@ Abstract
This study presents a novel framework to explore the fundamental relationship between gravity and quantum mechanics, with particular emphasis on the role of primordial black holes (PBHs) in cosmology. Through the concept of self-gravitating condensed light, specifically in the form of the experimentally discovered photon Bose-Einstein condensate, this work examines the quantum attributes of PBHs and their implications for early universe dynamics, baryogenesis, and the very early formation of galaxies and large-scale structures, established by JWST data. The model also investigates quantized characteristics of PBHs, such as mass, entropy, and temperature, suggesting that quantum processes are fundamental to black hole mechanics. By addressing issues like the cosmological constant problem and the information loss paradox, dark matter, and dark energy, this work provides insights into Planck-scale physics and proposes that PBHs may serve as a bridge between quantum theory and general relativity. This study ultimately posits that understanding PBH physics is essential to resolving major cosmological and astrophysical paradoxes for the ultimate unification of quantum mechanics with gravity.
@@ II The sketches of “Einstein’s algebra”
We begin this section with Albert Einstein’s final work in 1956 [31], where he articulated a concluding statement that reflects his core methodological principle: “… a finite system of finite energy can be completely described by a finite set of quantum numbers.” This does not seem to align with a continuum theory and must lead to an attempt to find a purely algebraic theory for describing reality. However, nobody knows how to establish the foundation for such a theory.
@@
Inspired by Einstein’s vision, we aim to develop a theory rooted purely in algebra and try to bridge the seemingly insurmountable gap between gravity and quantum mechanics.
@@
In 1985, Price and Thorne introduced the ”membrane formalism” for black holes [32], which reframes the black hole’s horizon as a ”stretched horizon”—a 2D membrane within 3D space that responds dynamically to external forces. This membrane, building on Damour and Znajek’s work [33, 34, 35], acts like a viscous fluid with electrical charge and conductivity, as well as finite entropy and temperature, though it cannot conduct heat. The membrane’s interactions with the universe are governed by familiar fluid dynamics and thermodynamic principles, such as the Navier-Stokes equation, Ohm’s law, and the laws of thermodynamics. These concepts provide astrophysicists with intuitive tools for understanding black hole behavior in complex environments. In this context, we discuss the following two works. The discussion covers two works focused on early universe phenomena. In Grasso et al. [36], the authors explore the possibility that galactic and cluster magnetic fields have a primordial origin. They suggest that searching for imprints of these magnetic fields in the temperature and polarization anisotropies of the cosmic microwave background radiation (CMBR) is the best approach for testing this hypothesis. Another work done by Bruce et al. [37] proposes that dark energy could originate from dark matter condensation at a low redshift. They analyze constraints from CMBR, large-scale structure, supernovae, and radio galaxies using a four-parameter likelihood analysis, factoring in the Hubble constant and three parameters of the condensate field, $Q$.
@@
The above-mentioned concepts may also apply to 2D photon Bose-Einstein condensates (BEC) with rest energy (arise due to condensation) in various phenomena [11, 13, 38, 39]. The idea of a 2D photon BEC with rest energy can be linked to the early universe, particularly concerning PBHs, analog gravity, and related topics [40, 41, 42, 43]. This correlation can be framed within a physical model where PBHs are represented as 2D spherical photon condensates contained within their own gravitational fields [44]. Such a model naturally leads to a geometrical equation that directly connects the Compton wavelength, $\lambda_{sm}$, of condensed light (quantum theory) with the geodesic length (general relativity) [44]:
@@
where $R_{s}$ is Schwarzschild radius. From the preceding considerations, we can intuitively say that spacetime instructs condensed light quanta on how to curve, while condensed light quanta instructs spacetime on how to quantize. The expression for Schwarzschild radius in term of mass ($M_{s}$) of the PBH, may be expressed as
@@
$I_{s}=N_{s}=2n_{s}(bit);\leavevmode\nobreak\ \tau_{s}=\frac{4\pi t_{p}n_{s}^{\frac{3}{2}}}{3};\leavevmode\nobreak\ L_{s}=\frac{L_{p}}{4\pi n_{s}},$ (7)
@@
III Planck equivalence and corresponding relations
@@
[…]
$=\frac{\hbar}{l_{p}}\leavevmode\nobreak\ (de\leavevmode\nobreak\ Broglie);$
@@ 2. Secondly, in our framework, the gravitational squeezing force, expressed as:
$=\frac{dE_{s}}{dR_{s}}=\frac{d}{dR_{s}}(\frac{M_{p}c^{2}R_{s}}{2l_{p}})=\frac{c^{4}}{2G}=\frac{F_{p}}{2}$ (10)
@@
$\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)$
$=\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)+\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)$ (15)
@@
$\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)$
$=2\times\frac{E_{p}}{4\sqrt{n_{s}}}\leavevmode\nobreak\ (two-particle\leavevmode\nobreak\ energy)$ (16)
@@
$\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)$
$=\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)-\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)$
@@
$\Delta S_{s,s-1}=\frac{\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)}{T_{s}\leavevmode\nobreak\ (rad)}$
$=\Delta S_{s,s-1}\leavevmode\nobreak\ qu(rad)+\Delta S_{s,s-1}\leavevmode\nobreak\ gr(rad)$ (18)
@@
$\Delta S_{s,s-1}\leavevmode\nobreak\ gr(rad)=2k_{B}\leavevmode\nobreak\ \big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)},$
@@
    $\Delta E_{s,s-1}(rad)$
    $=\frac{\hbar}{c}a_{s}+\frac{\hbar}{c}a_{s}\big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)}$ (21)
@@
    $T_{s}(rad)=\frac{E_{p}}{4k_{B}\sqrt{n_{s}}}$
    $=\frac{\hbar a_{p}}{4ck_{B}\sqrt{n_{s}}}=\frac{\hbar}{2ck_{B}}a_{s}=T_{s}.$ (22)
@@ V PBHs Lifetime and Luminosity: Addressing the Information Loss Paradox
$=\frac{4\pi G^{2}M_{s}^{3}}{3\hbar c^{4}}$
$=\frac{\tau_{s}\leavevmode\nobreak\ (Hawking)}{3840}$ (26)
@@
$=\frac{L_{p}}{4\pi n_{s}}=\frac{\hbar c^{6}}{4\pi G^{2}M_{s}^{2}}$
$=3840L_{s}\leavevmode\nobreak\ (Hawking),$ (27)
@@
$I_{s}(rad)=\frac{\sum_{n=1}^{n_{s}}\Delta S_{s,s-1}qu(rad)}{k_{B}}=2n_{s}=I_{s}$ (30)
@@ VI Quantum Absorption of Free Light: Origins of Cosmic Disks and Jets
Considering the quantum characteristics of condensed light in PBHs, it is evident that only free light photons with a wavelength not exceeding the geodesic length of the PBHs are subject to accretion by quantum absorption. In other words, PBHs consume only free light or condensed light (such as from other black holes through collisions and mergers). All ordinary matter is melted in accretion disks and expelled in jets, accompanied by swirling magnetic fields through “funnel” and “slingshot” effects. As will be shown below, this is precisely what can explain the exceptionally early formation of the SMBHs from a huge reservoir of free light energy radiated by PBHs.
@@ VII Conservation of Quantum Information and Gravitational Radiation from PBH merger
$E_{s1,2}(gr(rad))$
$=E_{p}(\sqrt{n_{s1}}+\sqrt{n_{s2}}-\sqrt{n_{s1}+n_{s2}}).$ (39)
@@ VIII Planck Scale to Expansion: The Quantum Foundations of Cosmology
    $E(BH_{p})=E(\gamma_{p},\gamma_{p})-E(\Gamma_{p})=2E_{p}-E_{p}=E_{p}.$
@@
$M_{s}=M_{p}\frac{\Omega_{L}+\Omega_{DM}}{\aleph}=\frac{1+5\aleph}{4\aleph}\approx 2.945498M_{p}$ (60)

### v3

# Connecting Gravity and Quantum Physics: Primordial Black Holes and Accelerated Evolution of the Universe

@@ Abstract
This study presents a new framework to explore the fundamental relationship between gravity and quantum mechanics, with particular emphasis on the fundamental role of primordial black holes (PBHs) in cosmology. Through the concept of self-gravitating condensed light in the form of the experimentally discovered quantum photon Bose-Einstein condensate, this work examines the quantized gravitational, informational, thermodynamical, traditional, and other attributes of PBHs and their implications for early universe dynamics, baryogenesis, the very early formation of galaxies, supermassive black holes (SMBHs) and large-scale structures. Precise calculations have shown that primordial protogalaxies with supermassive black holes at their centers of gravity were formed before the recombination epoch. By solving issues like the cosmological constant problem and the information loss paradox, dark matter and dark energy, this work provides insights into Planck-scale physics and it’s impact to cosmology. In such a way PBHs serve as a bridge between quantum theory and general relativity. This study ultimately posits that presented PBH physics is essential to resolving major cosmological and astrophysical issues, paradoxes and ”mysteries”, such as the accelerated evolution of the Universe established by JWST and other observations.
@@ I Introduction
The rapid advancement of modern experimental, observational, and measurement technologies—delivering precise and often unexpected data, as exemplified by the JWST—revives longstanding, unresolved questions in theoretical physics and cosmology. At the heart of these open problems lies the fundamental, yet elusive, connection between General Relativity and Quantum Theory.
The main goal of our work is to find and mathematically express this connection, and, on it’s basis, to resolve or remove such major issues, paradoxes and ”mysteries” as cosmological constant problem, informational loss paradox, cosmological principle problem related to the large-scale structure ”excesses”, physical nature and origin of dark matter and dark energy, baryogenesis, and baryonic assymetry, the very early growth of SMBHs as the gravitational centers of the very early proto-galaxies, the very early clumping of the Universe and it’s accelerated evolution from the Beginning of Time.
Since S. Hawking’s time more and more cosmologists have begun to recognize the exceptional role of primordial black holes in cosmogenesis and cosmic evolution, but no one knew what kind of gravitating matter they were made of and what their real physical nature is. It should also be noted that since the time of J.Beckenstein, V.Mukhanov, and G.t’Hooft many people have also begun to talk about the quantum nature of black holes, comparing it with the quantum nature of hydrogen atom.
Fortunately, the recent experimental discovery of quantized condensed light predicted a century ago by A. Einstein in the form of photon Bose-Einstein condensate (photon BEC) gives us a unique opportunity to relate its Compton wavelength to the gravitational radius of a primordial black hole. This geometrical approach makes PBHs a unique bridge between quantum theory and general relativity.
It is very important to note that photons are the only particles whose wavelength is capable to cover the immense range of black hole sizes, ranging from Planck-scale sizes to tens of solar system diameters. This is what makes condensed light with its effective rest energy and effective rest mass the only candidate for the ”substance” that black holes are made of.
It is the quantum nature of gravitating condensed light that predicts the quantized characteristics of the PBHs such as mass, amount of quantum information, entropy, temperature, radiation, quantum absorption, lifetime, etc. Also, it provides the simple and clear laws that govern their birth and death, their development, and their decisive role in the accelerated evolution of the Universe. Moreover, presented PBHs physics opens the door to the Terra incognita of Planck-scale physics and the beginning of time.
It should be especially noted that in this work we assume that modern dark matter consists of primordial black holes with a typical mass of 1.659x1018gr and a minimum of 2.598x1017gr, while dark energy is represented by the gravitational radiation associated with Dark Matter mass-energy loss processes, starting from the hot Big Bang to the present time.
Moreover, the precise calculations presented in this paper indicate that the first protogalaxies, harboring supermassive black holes and composed of both ordinary baryonic and dark matter, emerged before the epoch of recombination. At this stage, the Universe existed as a primordial, clumped structure—a ”universal supercluster.” As cosmic expansion progressed, this vast structure fragmented, giving rise to the immense supervoids, filaments, and superclusters that define the large-scale cosmic web we observe today.
Let us introduce a brief look at a number of theoretical and practical studies that are related to the topic of our research.
@@ II The sketches of “Einstein’s algebra”
We begin this section with Albert Einstein’s final work in 1956 [31], where he articulated a concluding statement that reflects his core methodological principle: ”… a finite system of finite energy can be completely described by a finite set of quantum numbers. This does not seem to align with a continuum theory and must lead to an attempt to find a purely algebraic theory for describing reality. However, nobody knows how to establish the foundation for such a theory”.
@@
Inspired by Einstein’s vision, we aim to develop a description of physical reality rooted purely in algebra and try to bridge the seemingly insurmountable gap between gravity and quantum mechanics.
As it was mentioned at the beginning of our Introduction, only the gravitating condensed light with its practically limitless range of Compton wavelengths from Planck-scale to tens of diameters of Solar System can serve as the ”building material” for black holes. Its quantum nature must be inevitably reflected in the quantized characteristics of black holes. But what is their geometrical structure? What is the constructive design of such unique ”buildings” of the observable Universe?
We find a leading answer in the well-established numerical relativity from the well-proven method of the ”simple excision of a black hole in 3+1 numerical relativity” [109], where black hole is presented as a very thin layer of gravitating matter, the internal spacetime of which is completely excluded from the calculation.
In this study, we are dealing with a extremely thin ”membrane” model of black holes. It is not difficult to imagine what would happen if we replace abstract matter with the experimentally discovered 2D quantum photon condensate predicted by the same A. Einstein a hundred years ago [13].
@@
Historically, in 1985, Price and Thorne introduced the ”membrane formalism” for black holes [32], which reframes the black hole’s horizon as a ”stretched horizon”—a 2D membrane within 3D space that responds dynamically to external forces. This membrane, building on Damour and Znajek’s work [33, 34, 35], acts like a viscous fluid with electrical charge and conductivity, as well as finite entropy and temperature, though it cannot conduct heat. The membrane’s interactions with the universe are governed by familiar fluid dynamics and thermodynamic principles, such as the Navier-Stokes equation, Ohm’s law, and the laws of thermodynamics. These concepts provide astrophysicists with intuitive tools for understanding black hole behavior in complex environments. In this context, we discuss the following two works. The discussion covers two works focused on early universe phenomena. In Grasso et al. [36], the authors explore the possibility that galactic and cluster magnetic fields have a primordial origin. They suggest that searching for imprints of these magnetic fields in the temperature and polarization anisotropies of the cosmic microwave background radiation (CMBR) is the best approach for testing this hypothesis. Another work done by Bruce et al. [37] proposes that dark energy could originate from dark matter condensation at a low redshift. They analyze constraints from CMBR, large-scale structures, supernovae, and radio galaxies using a four-parameter likelihood analysis, factoring in the Hubble constant and three parameters of the condensate field.
@@
The above-mentioned concepts, especially the concept of ”membrane paradigm”, may also apply to 2D photon Bose-Einstein condensates (BEC) with rest energy (arise due to condensation) in various phenomena [11, 13, 38, 39]. The idea of a 2D photon BEC with rest energy and rest mass can be linked to the early universe, particularly concerning PBHs, analog gravity, and related topics [40, 41, 42, 43].
The model of PBHs we present in this work is framed as 2D spherical photon condensates contained within their own gravitational fields [44]. Such a model naturally leads to a geometrical equation that directly connects the Compton wavelength, $\lambda_{sm}$, of condensed light (quantum theory) with the geodesic length (general relativity) [44]:
@@
where $R_{s}$ is Schwarzschild radius. To paraphrase John Wheeler’s famous quote, we can say that ”The Spacetime tells to condensed Light Quanta how to curve, the condensed Light Quanta tells to Spacetime how to quantize”. The expression for Schwarzschild radius in term of mass ($M_{s}$) of the PBH, may be expressed as
@@
$I_{s}=N_{s}=2n_{s}(bit);\leavevmode\nobreak\ \tau_{s}=\frac{4\pi t_{p}n_{s}^{\frac{3}{2}}}{3};\leavevmode\nobreak\ L_{s}=\frac{L_{p}}{4\pi n_{s}},$
@@
It should be noted that the quantization approach outlined for Schwarzschild black holes can be naturally extended to Kerr black holes.
The size of the equatorial circumference horizon of a Kerr black hole (the maximal geodesic length) is equal to [45]
$K(equatorial\leavevmode\nobreak\ circumference)=\frac{4\pi GM_{k}}{c^{2}}.$ (9)
where $M_{k}$ is Kerr black hole mass, $M_{s}$ is the mass of the corresponding Schwarzschild black hole taken as the irreducible one, $J$ is angular momentum. Note that when $J=0$ a Kerr black hole becomes the Schwarzschild one, with its proper form of the geodesic length:
$K(equatorial\leavevmode\nobreak\ circumference,J=0)=2\pi R_{s}.$ (10)
Putting the natural geometrical equation
$\lambda_{km}=\frac{4\pi GM_{k}}{c^{2}}$ (11)
and Compton wavelength definition
$\lambda_{km}=\frac{2\pi\hbar}{M_{km}c}$ (12)
we find
$n_{k}=\frac{M_{k}^{2}}{M_{p}^{2}}$ (13)
and other important formulas such as
$M_{k}=M_{p}\sqrt{n_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ I_{k}=N_{k}=2n_{k},\leavevmode\nobreak\ \leavevmode\nobreak$
$S_{k}=2k_{B}n_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ T_{k}=\frac{T_{p}}{4\sqrt{n_{k}}}etc.$ (14)
Starting from well known Kerr black hole formulas [45]:
1) Fundamental relation between $M_{k}$, $M_{s}$ and $J$
$M_{k}=\sqrt{M_{s}^{2}+\frac{J^{2}c^{2}}{4M_{s}^{2}G^{2}}},$ (15)
2) The fundamental condition for Kerr black hole outer horizon existence:
$J\leq\frac{GM_{k}^{2}}{c},$ (16)
we can quantize these relations by the natural substitution:
$M_{k}=M_{p}\sqrt{n_{k}},\leavevmode\nobreak\ \leavevmode\nobreak\ M_{s}=M_{p}\sqrt{n_{s}},$ (17)
and transform (15) and (16) as follows:
$n_{k}=n_{s}+\frac{J^{2}c^{2}}{4M_{p}^{4}G^{2}n_{s}},$ (18)
$J\leq\hbar n_{k}.$ (19)
Note that the right side of this fundamental inequality gives a very interesting physical meaning because we must take into account that the spin of condensed photons is equal to the reduced Planck constant and that the Kerr quantum number is equal to half of the number of condensed photons.
The above results concerning quantized Kerr black holes are of undoubted interest; however, we will not use them further, since the mechanisms of formation of angular momentum, $J$, manifest themselves only in the case of stellar mass and supermassive black holes.
III Planck equivalences and corresponding relations
@@
Planck gravity: $g_{p}=G\frac{M_{p}}{l_{p}^{2}}=\frac{F_{p}}{M_{p}}=a_{p}\leavevmode\nobreak$
@@ 2. Secondly, in our framework, the gravitational squeezing force, expressed as:
$=\frac{dE_{s}}{dR_{s}}=\frac{d}{dR_{s}}(\frac{M_{p}c^{2}R_{s}}{2l_{p}})$
$=\frac{c^{4}}{2G}=\frac{F_{p}}{2}\leavevmode\nobreak\ \leavevmode\nobreak$ (21)
@@
$\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)=\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)+\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)$
@@
$\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)=2\times\frac{E_{p}}{4\sqrt{n_{s}}}\leavevmode\nobreak\ (two-particle\leavevmode\nobreak\ energy)$
@@
$\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)=\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)-\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)$
@@
$\Delta S_{s,s-1}=\frac{\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)}{T_{s}\leavevmode\nobreak\ (rad)}=\Delta S_{s,s-1}\leavevmode\nobreak\ qu(rad)+\Delta S_{s,s-1}\leavevmode\nobreak\ gr(rad)$
@@
$\Delta S_{s,s-1}\leavevmode\nobreak\ gr(rad)=2k_{B}\leavevmode\nobreak\ \big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)},$ (30)
@@
    $\Delta E_{s,s-1}(rad)=\frac{\hbar}{c}a_{s}+\frac{\hbar}{c}a_{s}\big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)}$ (32)
@@
    $T_{s}(rad)=\frac{E_{p}}{4k_{B}\sqrt{n_{s}}}=\frac{\hbar a_{p}}{4ck_{B}\sqrt{n_{s}}}=\frac{\hbar}{2ck_{B}}a_{s}=T_{s}.$ (33)
@@ V PBHs Lifetime and Luminosity: Addressing the Information Loss Paradox
$=\frac{4\pi G^{2}M_{s}^{3}}{3\hbar c^{4}}=\frac{\tau_{s}\leavevmode\nobreak\ (Hawking)}{3840}$ (37)
@@
$=\frac{L_{p}}{4\pi n_{s}}=\frac{\hbar c^{6}}{4\pi G^{2}M_{s}^{2}}=3840L_{s}\leavevmode\nobreak\ (Hawking),$ (38)
@@
$I_{s}(rad)=\frac{\sum_{n=1}^{n_{s}}\Delta S_{s,s-1}qu(rad)}{k_{B}}=2n_{s}=I_{s}\leavevmode\nobreak\ \leavevmode\nobreak$ (41)
@@ VI Quantum Absorption of Free Light: Origins of Cosmic Disks and Jets
Considering the quantum characteristics of condensed light in PBHs, it is evident that only free light photons with a wavelength not exceeding the geodesic length of the PBHs are subject to accretion by quantum absorption. In other words, PBHs consume only free light or condensed light (such as from other black holes through collisions and mergers). All ordinary matter is melted in accretion disks and expelled in jets, accompanied by swirling magnetic fields through “funnel” and “slingshot” effects.
It is particularly noteworthy that this phenomenon precisely reflects the law of total baryon number conservation. A particular impressive manifestation of these effects, behind which this most important conservation law is hidden, are jets whose length is tens of times more than the size of the galaxies that it produced, such as M87 and Porphyrion. As will be shown below, this is precisely what can explain the exceptionally early formation of the SMBHs from a huge reservoir of free light energy radiated by PBHs.
@@ VII Conservation of Quantum Information and Gravitational Radiation from PBH merger
$E_{s1,2}(gr(rad))=E_{p}(\sqrt{n_{s1}}+\sqrt{n_{s2}}-\sqrt{n_{s1}+n_{s2}}).$
@@ VIII Planck Scale to Expansion: The Quantum Foundations of Cosmology
    $E(BH_{p})=E(\gamma_{p},\gamma_{p})-E(\Gamma_{p})=2E_{p}-E_{p}=E_{p}.$ (57)
@@
$M_{s}=M_{p}\frac{\Omega_{L}+\Omega_{DM}}{\aleph}=\frac{1+5\aleph}{4\aleph}\approx 2.945498M_{p}$
@@
Corollarium. This is an exceptionally important section of our research, which, based on the unified (quantum-mechanical and gravitational) algebraic description of reality bequeathed by A. Einstein, allows us not only to penetrate into unattainable Beginning of Time, but also to understand the great importance of ”quantum cosmology” and the accelerated evolution of the Universe. Let us list some important lessons of this research.
1. Moving backwards along the ”arrow of time” we find our observable Universe as a ”granular” spherical lump sized as an atom of gold, consisting of Planck 3D photon condensate, in a state with extremely high energy and mass density, temperature and pressure, as well extremely low entropy and lifetime, all expressed in Planck units.
[…]
4. It becomes clear that each dying generation of PBHs, through radiative transfer, ensures the growth of a new, less numerous generation of heavier ones, and that this process acquires a cyclical and synchronous character.